# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ 2.0.2. Example
Let $\mathcal{G}_{0},\dots,\mathcal{G}_{n}$ be any sequence as defined above such that $\mathcal{G}_{0}=\{G\cup f\}$. Lemma 3 shows that $(G,f)$ has the $d$-SIP if and only if, for each set $\mathcal{G}_{i}$ and each of its graphs $H$ that contain $f$, $(H\setminus f,f)$ has the $3$-SIP. This completes the proof. ∎
2.0.3. Example
Consider the graph-nonedge pairs $(G,f)$ (left) and $(G^{\prime},f)$ (right) in Figure 6, and let $f^{\prime}$ be the edge deleted from $G$ to obtain $G^{\prime}$. We saw in Example 1.2.1 that $(G,f)$ has the $2$-SIP while $(G^{\prime},f)$ does not. Since the endpoints of $f^{\prime}$ form a cs of $G\cup f$ and $G^{\prime}\cup f^{\prime}$, these facts can also be verified using Lemma 3. For any induced minor $[G]$, let $F^{\prime}$ be $\{[f]\}$ if $f$ is preserved in $[G]$, and the empty-set otherwise. Lemma 1 says that $([G],F^{\prime})$ has the $2$-SIP, which can be verified using Theorem 2. The contrapositive of Lemma 2 states that since $(G^{\prime},f)$ does not have the $2$-SIP, either $(G^{\prime}\cup f^{\prime},f)$ or $(G^{\prime},f^{\prime})$ does not have the $2$-SIP. Since some atom of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and is $K_{4}$, Theorem 2 shows that $(G^{\prime},f^{\prime})$ does not have the $2$-SIP, which verifies this statement. Lastly, observe that presence of $f^{\prime}$ separates $f$ from any atom with a $K_{4}$ minor, while the absence of $f^{\prime}$ causes $(G^{\prime},f)$ to inherit the non-$2$-SIP from $(G^{\prime},f^{\prime})$. Corollary 2 states that this situation can only occur if $G\cup f$ is not an atom.
2.1. Proof obstacles
Consider a graph-nonedge pair $(G,f)$. The right-hand side of Theorem 4 appears similar to a finite forbidden minor characterization, but has two key differences. First is that only the atoms of $G\cup f$ that contain $f$ must avoid $3$-forbidden minors, and second is that some such minor must preserve $f$. Using Lemma 3, we can choose $(G,f)$ so that it has the $3$-SIP and some atom of $G\cup f$ that does not contain $f$ has a $3$-forbidden minor. Furthermore, even if some atom of $G\cup f$ that contains $f$ has a $3$-forbidden minor, if no such minor preserves $f$, then $(G,f)$ may still have the $3$-SIP. We exemplify the intricacies of the theorem by examining how the characterization of graph-nonedge pairs with the $d$-SIP changes from $d=1$ to $d=3$. A pair of vertices $uv$ of $G$ is doubled in a minor $[G]$ if it is preserved and $[uv]^{-1}$ contains at least one edge if $uv$ is a nonedge, or at least two edges if $uv$ is an edge. Also, $uv$ is retained in $[G]$, in which case $[G]$ is $uv$-retaining, if it is preserved and not doubled. See Figure 7. For each integer $d\leq 3$, we will note whether edge deletions preserve the $d$-SIP and which of the following statements are equivalent to $(G,f)$ not having the $d$-SIP:
  1. (i)
    Some atom of $G\cup f$ that contains $f$ has a $d$-forbidden minor.
  2. (ii)
  3. (iii)
  4. (iv)
@@ 2.1. Proof obstacles
    $G\cup f$ has some $f$-preserving $d$-forbidden minor.
The $d=1$ case is well-behaved. Recall that $K_{3}$ is the only $1$-forbidden minor. Figure 1(a) shows the smallest graph-nonedge pair that does not have the $1$-SIP. This example evidences the following equivalence: $(G,f)$ does not have the $1$-SIP if and only if (v) some cycle in $G$ contains $f$. Using this fact, it is easy to see that the family $A^{d}$ of graph-nonedge pairs that have the $1$-SIP is closed under both edge deletions and contractions. The following cycle of implications shows that (v) is equivalent to (i) - (iv): (v) $\implies$ (i) $\implies$ (ii) $\implies$ (iii) $\implies$ (iv) $\implies$ (v). The implication (v) $\implies$ (i) follows easily from Lemma 66 in Appendix E, and the other implications have straightforward proofs.
Once we move to $d=2$, some of the above implications become false. Recall that $K_{4}$ is the only $2$-forbidden minor. Figure 1(c) shows the smallest graph-nonedge pair that does not have the $2$-SIP. Theorem 2 and Theorem 5.2 in [22] show that $(G,f)$ not having the $2$-SIP is equivalent to (i) - (iii). While (i), (ii), or (iii) $\implies$ (iv), the converse is false. For example, see the graph-nonedge pair on the left in Figure 6. Also unlike the $d=1$ case, the above equivalences can be used to show that $A^{d}$ is not closed under edge deletions. See Example 1.2.1.
[…]
To summarize, $(G,f)$ may have the $3$-SIP even though some atom of $G\cup f$ that contains $f$ has a $3$-forbidden minor, so long as $f$ is contracted in any such minor. On the other hand, $(G,f)$ may not have the $3$-SIP even though $f$ cannot be retained in any $3$-forbidden minor. The latter statement results in infinitely many minimal graph-nonedge pairs (Definition 3), with respect to contractions, that do not have the $3$-SIP. This statement is false for any integer $d\leq 2$, and so there are finitely many minimal pairs that do not have the $d$-SIP. Furthermore, the $3$-SIP behaves badly when edges are deleted. For any graph-nonedge pair $(G^{\prime},f^{\prime})$ such that $G^{\prime}$ is obtained from $G$ by deleting some edge, it may be that $(G,f)$ has the $3$-SIP while $(G^{\prime},f^{\prime})$ does not, both $(G,f)$ and $(G^{\prime},f^{\prime})$ have the $3$-SIP, $(G,f)$ does not have the $3$-SIP while $(G^{\prime},f^{\prime})$ does, or both $(G,f)$ and $(G^{\prime},f^{\prime})$ do not have the $3$-SIP.
3. Proof sketch for the converse direction of Theorem 4
From here on, when we refer to a forbidden minor we mean a $3$-forbidden minor. Given a graph-nonedge pair $(G_{\star},f_{\star})$, we show that if no atom of $G_{\star}\cup f_{\star}$ that contains $f_{\star}$ has an $f_{\star}$-preserving forbidden minor, then $(G_{\star},f_{\star})$ has the $3$-SIP. The proof is by induction on the number $n$ of vertices in $G_{\star}$. When $n\leq 5$, Theorem 3 shows that $(G_{\star},f_{\star})$ has the $3$-SIP. In the inductive step, this theorem along with Corollary 3 allow us to assume that $G_{\star}\cup f_{\star}$ is an atom that has a forbidden minor. For the remainder of the argument, refer to Figure 11. Given a graph $G$ and one of its edges $e$, a minor $[G]_{A}$ is an $e$-winged graph minor if it is one of the graphs in Figure 8 such that $[e]$ is $w_{1}w_{2}$. Since atoms are $2$-connected, Proposition 1, stated below and proved in Appendix B.1, shows that $G_{\star}\cup f_{\star}$ has an $f_{\star}$-winged graph minor. See Figure 11(a).
Figure 11. See the discussion below. (A) an $f_{\star}$-winged graph minor $[G_{\star}\cup f_{\star}]$. (B) $[G_{\star}\cup f_{\star}]$ has singleton wing tips. (C) pins $\{x,y\}$ and $\{w,z\}$ of $[G_{\star}\cup f_{\star}]$. (D) $[G_{\star}\cup f_{\star}]$ is chained and $(G_{\star},\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected.
@@ 3. Proof sketch for the converse direction of Theorem 4
If $G$ is a $2$-connected graph and $e$ is one of its edges such that $G$ has a forbidden minor but no $e$-preserving forbidden minor, then $G$ has an $e$-winged graph minor.

### v2

@@ 4   Proof of the converse direction of the main Theorem 4
Let $e=uv$ and $[G]$ be a minor that is the graph in Figure 14(a) (resp. Figure 14(b)) with $[u]$ as $w_{1}$ and $[v]$ as $w_{2}$. Additionally, let $[u]^{-1}$ have minimum size over all such minors. Let $H$ be the subgraph of $G$ induced by $[u]^{-1}$ and $x$ be any vertex in $[u]^{-1}\setminus\{u\}$ with a neighbor in $[v]^{-1}$, which exists since $G$ is $2$-connected. Also, let $J$ be the connected component of $H\setminus x$ that contains $u$. First, we show how to choose $x$ so that $J$ contains some vertex with a neighbor in $G\setminus([u]^{-1}\cup[v]^{-1})$. If $J$ contains such a vertex, then our current choice for $x$ works. Otherwise, $G$ being $2$-connected gives a path in $G\setminus x$ between $u$ and some vertex in $G\setminus([u]^{-1}\cup[v]^{-1})$. Hence, some connected component $K$ of $H\setminus x$ contains both a vertex $y$ with a neighbor in $[v]^{-1}$ and a vertex with a neighbor in $G\setminus([u]^{-1}\cup[v]^{-1})$. Note that the connected component of $H\setminus y$ that contains $u$ has size strictly greater than $J$, since it contains both $J$ and $x$. Hence, we can set $x$ to be $y$ and repeat this argument until $J$ has the desired property. Consider the minor $[G]_{1}$ obtained via the vertex exchange from $[u]^{-1}$ to $[v]^{-1}$ that exchanges $x$ and fixes $J$. This operation along with our minimality assumption on $[u]^{-1}$ and our choice for $x$ ensure that $[G]_{1}$ has the desired properties. ∎
Figure 15: Graphs in Lemmas 4 and 5. Line-segments between $w_{2}$ and the blue circle represent an edge between $w_{2}$ and some vertex in the circle, but note that dashed line-segments may be nonedges. See also the proofs of Proposition 1 in this section, Lemma 9 in Section A.1, Lemma 29 in Section B.2, and Lemma 57 in Section B.3.
@@
If some minor $[G]$ of a graph $G$ is one of the graphs in Figure 15, then, for any edge $e$ in $[w_{1}w_{2}]^{-1}$, $G$ has an $e$-preserving forbidden minor.
Proof.
For each graph in Figure 15, a solid line-segment represents an edge and a dashed line-segment represents a nonedge. If a line-segment is between $w_{2}$ and the blue circle, then the endpoints of this edge or nonedge is $w_{2}$ and some vertex in the circle. The lemma follows from the observation that each of these graphs has a $w_{1}w_{2}$-preserving forbidden minor. ∎
@@
If $G$ is a $2$-connected graph with an edge $e$ and a forbidden minor but no $e$-preserving forbidden minor, then $G$ has an $e$-winged graph minor.
Proof.
Assume $G$ is $2$-connected and has a forbidden minor but no $e$-preserving forbidden minor. Then, $[e]$ is an edge of some minor $[G]$ whose contraction yields a forbidden minor. Hence, $[G]$ is one of the graphs in Figures 8, 15, or 14 with $[e]$ as $w_{1}w_{2}$, and thus applying Lemmas 4 and 5 completes the proof. ∎